# Gaussian Process Priors for Boundary Value Problems of Linear Partial Differential Equations

### v1

@@ Appendix C Ehrenpreis–Palamodov Theorem and EPGP
Next, we clarify the notion of smooth solution and the topology with respect to which we have density. First, we write $\mathcal{F}=C^{\infty}(\Omega)$ to be the space of smooth functions $\Omega\to\mathbb{C}$. This is a Frechét space under the standard topology induced by the semi-norms $s_{a,b}(u)=\max_{|\alpha|=a,x\in\Omega_{b}}|\partial^{\alpha}u(x)|$, where $\Omega_{b}\uparrow\Omega$ is an increasing sequence of compact sets which exhausts $\Omega$. This is to say that a sequence $u_{q}\to u$ in $\mathcal{F}$ if $s_{a,b}(u_{q}-u)\to 0$ for all $a,b$. Algebraically, $\mathcal{F}^{k}$ is an $R$-module under the action of differentiation.
@@
Our earlier algorithm EPGP from (Harkonen et al., 2023) revolves around fitting coefficients $c_{i,j,h}$ and “frequencies” $z_{i,j,h}$ in formula (EP). To simplify notation, we will simply write $\{b(x;z)\}_{z\in V}$ for the continuously indexed basis that we are working with (exponential-polynomial solutions of $A(\partial)u=0$). We will write our predictions in form
@@
Appendix D Free wave equation in 2D and 3D
In fact, our first improvement of EPGP is to update it to include initial speeds as well as initial conditions. We will explain this for the example of the wave equation in arbitrary dimension $n$.
Even without boundary conditions, this is an important example for which ongoing research is being developed (Henderson et al., 2023). Writing $\square u=u_{tt}-u_{x_{1}x_{1}}-u_{x_{2}x_{2}}-\ldots-u_{x_{n}x_{n}}$, the problem to consider is the Free Wave Equation:
@@ Appendix D Free wave equation in 2D and 3D
As an example, we will consider the two dimensional case,
@@
where $f(x)=\exp(-5x^{2})$ or $f(x)=\cos(5x)$. The true solution is given by $u(t,x,y)=f(x+t-2)+f(y+t-2)$ and our results can be found in Figure 9.
@@
Figure 9: Solutions to the initial boundary value problem for the 2-D wave equation (no boundary conditions). Solution fits both initial condition and initial speed. This is a necessary improvement of the EPGP algorithm from (Harkonen et al., 2023), which can produce non-physical solutions in certain cases. Animations can be found in the ancillary files exp free.mp4 and cos free.mp4.
Appendix E Check using conservation of energy
@@
[…]
@@
[…]
@@
Appendix F Wave equation in bounded domains
In this section we will provide numerical results for the $2$-D wave equation in bounded domains, meaning that we investigate $u_{tt}-(u_{xx}+u_{yy})=0\text{ in }\Omega\times(0,T)$ with Dirichlet or Neumann boundary conditions on $\partial\Omega$, where $\Omega\subset\operatorname*{\mathbb{R}}^{2}$ will is a bounded open convex set. We will use both EPGP (Section 3.1) and B-EPGP methods and compare them.
We will check the validity of our results using the conservation of energy principle from Theorem E.1. We will show that by using EPGP a non-negligible amount of energy is lost/dissipated. In fact, we can say more: We will solve initial boundary value problems with given initial condition and zero initial speed, meaning that we will solve
which is a well-posed problem under Dirichlet or Neumann boundary conditions. By Theorem E.1 we have
Thus in our experiments we will compare the energy of the B-EPGP with the true value computed from initial conditions above and also show its superiority to EPGP.
F.1 Rectangular domains
$\begin{cases}u_{tt}-(u_{xx}+u_{yy})=0&\text{for }x,y\in(0,4),\,t\in(0,12)\\ u(0,x,y)=\exp(-10((x-1)^{2}+(y-1)^{2}))&\text{for }x,y\in(0,4)\\ u_{t}(0,x,y)=0&\text{for }x,y\in(0,4)\\ u(t,x,y)=0&\text{for $x$ or $y=0$ or $4$.}\end{cases}$
We will give a shortcut to our B-EPGP algorithm for finding a basis. We consider $H_{1}=\{x=0\}$, $H_{2}=\{y=0\}$, $H_{3}=\{x=4\}$, $H_{4}=\{y=4\}$ and let $e^{at+bx+cy}$ be a solution of the wave equation, meaning that $a^{2}=b^{2}+c^{2}$. We can calculate a basis for the wedge $\{x,y>0\}$, see Appendix G.2, which is
$e^{at+bx+cy}-e^{at-bx+cy}-e^{at+bx-cy}+e^{at-bx-cy}$ $=e^{at}(e^{bx+cy}-e^{-bx+cy}-e^{bx-cy}+e^{-bx-cy})$
@@ F.1 Rectangular domains
$=2\sqrt{-1}e^{at}\sin(bx)(e^{cy}-e^{-cy})=-4e^{at}\sin(bx)\sin(cy)$
Here we observe a shortcut: if we set $b,c\in\mathbb{Z}$, we obtain a set of functions $\{e^{\pm\tfrac{\pi\sqrt{-1}}{4}\sqrt{j^{2}+k^{2}}t}\sin(\tfrac{\pi}{4}jx)\sin(\tfrac{\pi}{4}ky)\}_{j,k\in\mathbb{Z}}$ which satisfy the boundary condition on $H_{2}$ as well (see also Appendix G.1). That set of functions has linear span which is dense in the set of all solutions to (12) by Theorem B.1.
@@
Figure 10: Solution of 2-D wave equation in a rectangular domain calculated using the EPGP and B-EPGP methods. We use a Dirichlet boundary condition which is visible above from the fact that the edges of our plots have the same color in all snapshots. Animations can be found in the ancillary files rec-S.mp4 and rec-R.mp4.
Figure 11: Proof of energy conservation for the 2-D wave equation in a rectangular domain. EPGP method incurs a non-negligible loss of energy. The error in B-EPGP is partly due to our approximation of the energy integral in (11).
F.2 Triangular domains
@@
We will use the B-EPGP basis computed in Section F.1 and an odd reflection in the diagonal line $y=x$ which gives
@@
Figure 12: Solution of 2-D wave equation in a triangular domain with Dirichlet boundary conditions calculated using the EPGP and B-EPGP methods. Animations can be found in the ancillary files triangle.mp4 and triangleEPGP.mp4.
Figure 13: Proof of energy conservation for the 2-D wave equation in a triangular domain. S-EPGP method incurs a non-negligible loss of energy. The error in B-EPGP is partly due to our approximation of the energy integral in (11).
F.3 Circle: Drum membrane
Our methods extend to non polygonal domains. Here we use the direct method from Section 3.1 to cover the 2-D wave equation in a disc, a classic model for circular drum membranes. We will consider $\Omega=\{(x,y):x^{2}+y^{2}\leq 16\}$ and $t\in(0,7)$, and the equations are
@@ F.3 Circle: Drum membrane
Figure 14: Radially symmetric solution to 2-D wave equation in a circular domain with Dirichlet boundary conditions evaluated at 8 timepoints. The animation can be found in the ancillary file circle.mp4.
Figure 15: Proof of energy conservation for the 2-D wave equation in a circular domain. The error is partly due to our numerical approximation of the energy integral in (11).
Appendix G B-EPGP bases for wave equation in intersections of two halfspaces
We will also analyze the next simplest case which is more general than halfspaces (Appendix A), namely intersections of two halfspaces. These, have only two relative positions, parallel or not. When the boundary hyperplanes intersect, we will distinguish between the case when the angle is acute or not.
G.1 Parallel halfspaces: slabs
Consider the equation in Example 3.7, namely 1-D wave equation in an interval:
@@ G.1 Parallel halfspaces: slabs
We will show how to use our B-EPGP algorithm to calculate a basis. We consider $H_{1}=\{x=0\}$, $H_{2}=\{x=\pi\}$ and let $e^{at+bx}$ be a solution of the wave equation, meaning that $a^{2}=b^{2}$, so we will simply write $a=\pm b$. We obtain $e^{b(\pm t+x)}$. Using Section A, this basis is extended to a basis which satisfies the $H_{1}$ condition by
@@
Here we observe a shortcut: if we set $b\in\sqrt{-1}\mathbb{Z}$, we obtain a set of functions $\{e^{\pm\sqrt{-1}jt}\sin(jx)\}_{j\in\mathbb{Z}}$ which satisfy the boundary condition on $H_{2}$ as well. This set has linear span which is dense in the set of all solutions to (12) by Theorem B.1.
@@
Figure 16: B-EPGP solution for 1-D wave in an infinite slab with different boundary conditions. The difference color is explained by the different reflections: For Dirichlet, a negative wave is reflected (“hard boundary”) and for Neumann a positive wave is reflected (“soft boundary”). This behavior is already readable from the bases calculated in Appendix G.1.
G.2 Large wedges
@@
In Step 1 we notice that the boundary hyperplanes are $H_{1}=\{x=0\}$ and $H_{2}=\{y=0\}$.
In Step 2, we begin with one exponential solution
@@
This follows from the calculations in Section A.2. Similarly, we extend to a basis that satisfies the $H_{2}$ condition
@@
This gives us the intermediate “basis”, at the end of Step 2,
$b(z)=\{e^{at+bx+cy}-e^{at-bx+cy},e^{at+bx+cy}+e^{at+bx-cy}\colon a^{2}+b^{2}=c^{2}\,\}.$
We proceed with Step 3 and check the both boundary conditions. We notice that each type of basis element satisfies exactly one of the two boundary conditions, so we must return to Step 2.
@@
G.3 Small wedges
We will also consider and also implement the case of an acute wedge, e.g. $\Omega=\{(x,y):x>0,y<x\}$ and $t\in(0,8)$. We will look at the 2-D wave equation with Neumann boundary conditions
@@ G.3 Small wedges
Figure 17: Solution of 2-D wave equation in a $45^{\circ}$ wedge domain evaluated at 9 timepoints. Since the domain is unbounded, the wave leaves the domain in finite time. The animation can be found in the ancillary file goodwedge.mp4.
Appendix H Heat equation in 2-D
In this case, we can use the calculations in Section A.3 and G.2 to obtain the B-EPGP basis
for which we obtain the B-EPGP basis

### v2

@@ Appendix E Ehrenpreis–Palamodov Theorem and EPGP
Next, we clarify the notion of smooth solution and the topology with respect to which we have density. First, we write $\mathcal{F}=C^{\infty}(\Omega)$ to be the space of smooth functions $\Omega\to\mathbb{C}$. This is a Frechét space under the standard topology induced by the semi-norms $s_{a,b}(u)=\max_{|\alpha|=a,x\in\Omega_{b}}|\partial^{\alpha}u(x)|$, where $\Omega_{b}\uparrow\Omega$ is an increasing sequence of compact sets which exhausts $\Omega$. This is to say that a sequence $u_{q}\to u$ in $\mathcal{F}$ if $s_{a,b}(u_{q}-u)\to 0$ for all $a,b$. Algebraically, $\mathcal{F}^{k}$ is a $\mathbb{C}[\partial]$-module under the action of differentiation.
@@
The algorithm EPGP from [32] revolves around fitting coefficients $c_{i,j,h}$ and “frequencies” $z_{i,j,h}$ in formula (EP). To simplify notation, we will simply write $\{b(x;z)\}_{z\in V}$ for the continuously indexed basis that we are working with (exponential-polynomial solutions of $A(\partial)u=0$). Predictions are written in the form
@@
We use the same regression model in the present paper, by choosing $b(x;z)$ according to the B-EPGP base instead of the EPGP base.
Appendix F Free wave equation in 2D and 3D
In fact, our first improvement of EPGP is to update it to include initial velocity as well as initial displacement. We will explain this for the example of the wave equation in arbitrary dimension $n$.
Even without boundary conditions, this is an important example for which ongoing research is being developed [36]. Writing $\square u=u_{tt}-u_{x_{1}x_{1}}-u_{x_{2}x_{2}}-\ldots-u_{x_{n}x_{n}}$, the problem to consider is the Free Wave Equation:
@@ Appendix F Free wave equation in 2D and 3D
As an example, we will consider the 2D case,
@@
where $f(x)=\exp(-5x^{2})$ or $f(x)=\cos(5x)$. The true solution is given by u(t,x,y)=f(x+t2)+f(y+t2)u(t,x,y)=f(x+t-2)+f(y+t-2) and our results can be found in Figure 5.
@@
Figure 5: Solutions to the initial boundary value problem for the 2D wave equation (no boundary conditions). Solution fits both initial condition and initial speed. This is a necessary improvement of the EPGP algorithm from [32], which can produce non-physical solutions in certain cases. Animations can be found in the supplementary material as free_velocity1.mp4 and free_velocity2.mp4.
Appendix G Check using conservation of energy
@@
$E(t)=\int_{\Omega}\left|\frac{\partial u(x,t)}{\partial t}\right|^{2}+\sum_{j=1}^{d}\int_{\Omega}\left|\frac{\partial u(x,t)}{\partial x_{j}}\right|^{2}dx$
@@
$E^{\prime}(t)$ $=\int_{\Omega}\left[2u_{t}u_{tt}+2(\nabla u\cdot\nabla u_{t})\right]dx$
$=\int_{\Omega}2u_{t}u_{tt}dx-2\int_{\Omega}(\Delta u)u_{t}dx+2\int_{\partial\Omega}u_{t}\frac{\partial u}{\partial n}dS$
$=\int_{\Omega}2u_{t}(u_{tt}-\Delta u)dx+2\int_{\partial\Omega}u_{t}\frac{\partial u}{\partial n}dS$
$=2\int_{\partial\Omega}u_{t}\frac{\partial u}{\partial n}dS.$
@@
where $T$ is the final time ($t\in[0,T]$) and we choose $h=0.1$.
Appendix H Wave equation in bounded domains
In this section we will provide numerical results for the $2$-D wave equation in bounded domains, meaning that we investigate $u_{tt}-(u_{xx}+u_{yy})=0\text{ in }\Omega\times(0,T)$ with Dirichlet or Neumann boundary conditions on $\partial\Omega$, where $\Omega\subset\operatorname*{\mathbb{R}}^{2}$ will is a bounded open convex set. We will use both EPGP (Remark 3.1) and B-EPGP methods and compare them.
We will check the validity of our results using the conservation of energy principle from Theorem G.1. We will show that by using EPGP a non-negligible amount of energy is lost/dissipated. In fact, we can say more: We will solve initial boundary value problems with given initial condition and zero initial speed, meaning that we will solve
which is a well-posed problem under Dirichlet or Neumann boundary conditions. By Theorem G.1 we have
Thus in our experiments we will compare the energy of the B-EPGP with the true value computed from initial conditions above and also show its superiority over EPGP.
H.1 Rectangular domains
@@
We will give a shortcut to our B-EPGP algorithm for finding a basis. We consider $H_{1}=\{x=0\}$, $H_{2}=\{y=0\}$, $H_{3}=\{x=4\}$, $H_{4}=\{y=4\}$ and let $e^{at+bx+cy}$ be a solution of the wave equation, meaning that $a^{2}=b^{2}+c^{2}$. We can calculate a basis for the wedge $\{x,y>0\}$, see Appendix J.2, which is
$\phantom{=}e^{at+bx+cy}-e^{at-bx+cy}-e^{at+bx-cy}+e^{at-bx-cy}$
$=e^{at}(e^{bx+cy}-e^{-bx+cy}-e^{bx-cy}+e^{-bx-cy})$
@@
$=2\sqrt{-1}e^{at}\sin(bx)(e^{cy}-e^{-cy})$
$=-4e^{at}\sin(bx)\sin(cy)$
Here we observe a shortcut: if we set $b,c\in\mathbb{Z}$, we obtain a set of functions
$\{e^{\pm\tfrac{\pi\sqrt{-1}}{4}\sqrt{j^{2}+k^{2}}t}\sin(\tfrac{\pi}{4}jx)\sin(\tfrac{\pi}{4}ky)\}_{j,k\in\mathbb{Z}}$
which satisfy the boundary condition on $H_{2}$ as well (see also Appendix J.1). That set of functions has linear span which is dense in the set of all solutions to (12) by Theorem D.1.
@@
Figure 6: Solution of 2D wave equation in a rectangular domain calculated using the EPGP and B-EPGP methods. We use a Dirichlet boundary condition which is visible above from the fact that the edges of our plots have the same color in all snapshots. The results of B-EPGP seem more reasonable, as they are more stable and lack corrupting high frequencies. Animations can be found in the supplementary material as rectangle_EPGP.mp4 and rectangle_B-EPGP.mp4.
Figure 7: This figure shows the energy conservation for the 2D wave equation in a rectangular domain. We expect a constant value from physical principles. EPGP method incurs a non-negligible loss of energy. The error in B-EPGP is partly due to our approximation of the energy integral in (11).
H.2 Triangular domains
@@
We will use the B-EPGP basis computed in Section H.1 and an odd reflection in the diagonal line $y=x$ which gives
@@
Figure 8: Solution of 2D wave equation in a triangular domain with Dirichlet boundary conditions calculated using the EPGP and B-EPGP methods. Animations can be found in the supplementary material as triangle_EPGP.mp4 and triangle_B-EPGP.mp4.
Figure 9: This figure shows the of energy conservation for the 2D wave equation in a triangular domain. We expect a constant value from physical principles. S-EPGP method incurs a non-negligible loss of energy. The error in B-EPGP is partly due to our approximation of the energy integral in (11).
H.3 Circle: Drum membrane
Our methods extend to non polygonal domains. Here we consider the 2D wave equation in a disc with Dirichlet boundary conditions, a classic model for circular drum membranes. We will consider the spatial domain $\Omega=\{(x,y):x^{2}+y^{2}\leq 16\}$ and $t\in(0,7)$, and the equations
@@ H.3 Circle: Drum membrane
Here we cannot use B-EPGP for polygons or Hybrid B-EPGP since no piece of the boundary of $\Omega$ is polygonal. Instead, we will use the following implementation of the baseline method EPGP as described in Remark 3.1: We consider the EPGP basis for the 2D wave equation without boundary conditions $\{e^{\pm\sqrt{a^{2}+b^{2}}t+ax+by}\}_{a,b\in\mathbb{C}}$ and model the Dirichlet boundary condition as data $u(x_{h},y_{h},t_{h})=0$ for points $(x_{h},y_{h},t_{h})\in\partial\Omega\times(0,7)$. We also give data points to represent the initial conditions at $t=0$ as in the rest of the paper.
@@
Figure 10: Radially symmetric solution to 2D wave equation in a circular domain with Dirichlet boundary conditions evaluated at 8 timepoints. The animation can be found in the supplementary material as circle_EPGP.mp4.
Figure 11: This figure shows the energy conservation for the 2D wave equation in a circular domain. We expect a constant value from physical principles. The error is partly due to our numerical approximation of the energy integral in (11).
Appendix I B-EPGP and EPGP on 2D Wave Equation
Table 2: The results from Section I show the superiority of B-EPGP over EPGP when comparing the median L1-difference to the exact solution and the computation time. We report the standard deviation of ten repetitions.
Algorithm Abs Err($10^{-4}$) Rel Err(%) Time(s)
EPGP $4.26\pm 0.21$ $1.82\pm 0.09$ $6059$
B-EPGP (ours) $\textbf{0.48}\pm 0.02$ $\textbf{0.72}\pm 0.03$ $804$
We compare B-EPGP and EPGP in-depth, discussing accuracy and computational resources. Consider the following initial boundary value problem for the 2D wave equation in a halfspace $x>0$:
$\begin{cases}u_{tt}=u_{xx}+u_{yy}&\text{for $t\in(0,\infty),x\in(0,\infty),y\in\operatorname*{\mathbb{R}}$}\\ u(0,x,y)=f(x,y)\text{ and }u_{t}(0,x,y)=0&\text{for $x\in[0,\infty),y\in\operatorname*{\mathbb{R}}$}\\ u_{x}(t,0,y)=0&\text{for $t\in(0,\infty),y\in\operatorname*{\mathbb{R}}$}\end{cases}$
with initial condition $f=f_{1}+f_{2}+f_{3}$ for $c_{i}=-5i^{2}+20i-10$ and
$f_{i}(x,y)=$ $J_{0}(c_{i}\cdot\sqrt{((x-i)^{2}+(y-i)^{2})})+J_{0}(c_{i}\cdot\sqrt{((x+i)^{2}+(y-i)^{2})}),$
where $J_{0}$ is the Bessel function of order 0. The unique exact solution is given by
$u(t,x,y)$ $=f_{1}(x,y)\cos(5t)+f_{2}(x,y)\cos(10t)+f_{3}(x,y)\cos(5t).$
We can use B-EPGP to obtain the basis
$e^{\alpha x+\beta y+\tau t}+e^{-\alpha x+\beta y+\tau t}+e^{\alpha x+\beta y-\tau t}+e^{-\alpha x+\beta y-\tau t}\text{ for $\alpha,\beta,\tau\in\mathbb{C}$ and $\alpha^{2}+\beta^{2}=\tau^{2}$.}$
We fix the number of frequencies to be the same in both EPGP and B-EPGP. Hence, B-EPGP requires only one-quarter as many trainable parameters as EPGP. Both methods model initial data on a spatial grid of spacing 0.2. EPGP requires additional boundary data sampled on a grid, which triples the number of data and increases the size of the covariance matrix by a factor of nine, leading to substantially higher computational cost. A one-sided paired Wilcoxon signed-rank test over ten repetitions, with the alternative hypothesis that the location shift is less than zero, yields a highly significant $p$-value of $0.001\ll 0.05$. Table 2 sums up the comparison, where B-EPGP reduces prediction error, runtime, and memory usage by an order of magnitude. Figure 13 describes how EPGP and B-EPGP perform with different data size and basis elements. A visual comparison of the two solutions can be seen in Figure 12.
(a) $t=0.0$,
EPGP
(b) $t=0.5$,
EPGP
(c) $t=1.0$,
EPGP
(d) $t=1.5$,
EPGP
(e) $t=2.0$,
EPGP
(f) $t=2.5$,
EPGP
(g) $t=3.0$,
EPGP
(h) $t=3.5$,
EPGP
(i) $t=4.0$,
EPGP
(j) $t=0.0,$
B-EPGP
(k) $t=0.5,$
B-EPGP
(l) $t=1.0,$
B-EPGP
(m) $t=1.5,$
B-EPGP
(n) $t=2.0,$
B-EPGP
(o) $t=2.5,$
B-EPGP
(p) $t=3.0,$
B-EPGP
(q) $t=3.5,$
B-EPGP
(r) $t=4.0,$
B-EPGP
Figure 12: Predictions to the 2D wave equation from Section I using EPGP and B-EPGP. We use this experiment as a benchmark as we can compare both predictions with a highly nontrivial exact solution. Animations can be found in the supplementary material as halfplane_EPGP.mp4 and halfplane_B-EPGP.mp4.
(a) $t=0.0$
(b) $t=0.1$
(c) $t=0.2$
(d) $t=0.3$
Figure 13: The relationship between the error and the data size / number of basis elements we use, for both EPGP and B-EPGP, and both in normal scale and log scale
Appendix J B-EPGP bases for wave equation in intersections of two halfspaces
We will also analyze the next simplest case which is more general than halfspaces (Appendix C), namely intersections of two halfspaces. These, have only two relative positions, parallel or not, in which case we will distinguish between the case when the angle is acute or not.
J.1 Parallel halfspaces: slabs
Consider the equation in Example 3.8, namely 1D wave equation in an interval:
@@ Appendix I B-EPGP and EPGP on 2D Wave Equation
We will show how to use our B-EPGP algorithm to calculate a basis. We consider $H_{1}=\{x=0\}$, $H_{2}=\{x=\pi\}$ and let $e^{at+bx}$ be a solution of the wave equation, meaning that $a^{2}=b^{2}$, so we will simply write $a=\pm b$. We obtain $e^{b(\pm t+x)}$. Using Section C, this basis is extended to a basis which satisfies the $H_{1}$ condition by
@@
Here we observe a shortcut: if we set $b\in\sqrt{-1}\mathbb{Z}$, we obtain a set of functions $\{e^{\pm\sqrt{-1}jt}\sin(jx)\}_{j\in\mathbb{Z}}$ which satisfy the boundary condition on $H_{2}$ as well. This set has linear span which is dense in the set of all solutions to (12) by Theorem D.1.
@@
Figure 14: B-EPGP solution for 1D wave in an infinite slab with different boundary conditions. The difference color is explained by the different reflections: For Dirichlet, a negative wave is reflected (“hard boundary”) and for Neumann a positive wave is reflected (“soft boundary”). This behavior is already readable from the bases calculated in Appendix J.1.
J.2 Large wedges
@@
We notice that the boundary hyperplanes are $H_{1}=\{x=0\}$ and $H_{2}=\{y=0\}$.
In Step (i), we begin with one exponential solution
@@
This follows from the calculations in Section C.2. Similarly, we extend to a basis that satisfies the $H_{2}$ condition
@@
This gives us the intermediate “basis”, at the end of Step (ii),
$\{e^{at+bx+cy}-e^{at-bx+cy},e^{at+bx+cy}+e^{at+bx-cy}\colon a^{2}+b^{2}=c^{2}\,\}.$
We proceed with Step (i’) and check the both boundary conditions. We notice that each type of basis element satisfies exactly one of the two boundary conditions, so we must return to Step (ii).
@@
J.3 Small wedges
We will also consider and also implement the case of an acute wedge, e.g. $\Omega=\{(x,y):x>0,y<x\}$ and $t\in(0,8)$. We will look at the 2D wave equation with Neumann boundary conditions
@@
Figure 15: Solution of 2D wave equation in a $45^{\circ}$ wedge domain evaluated at 9 timepoints. Since the domain is unbounded, the wave leaves the domain in finite time. The animation can be found in the suplementary material as wedge_B-EPGP.mp4.
Appendix K Heat equation in 2D